• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:59:40 on December18,2006

    Entry number 190952

    keyword=shift summary

    Day Shift Summary:

    Shift workers: J. P. Chen (SL), T. Holmstrom (TO).

    8:00 data taking 90 uA on He3 target for kin. # 8.
    10:00 stopped beam for Mott measurement.
    10:35 moved left spectrometer to 30.2 degrees.
    set left spectrometer momentum to 2.227 GeV/c for quasi-elastic
    kinematics for target density current scan.
    10:50 All Quads were set. Left dipole no resposne, reboot ioc14.
    Still no response.
    11:00 Called tech-on-call. Todd and Heidi came to check.
    11:15 requested controlled access.
    12:10 Techs came out of the hall. Dipole is ramping back up.
    Hall back to power permit (later beam permit).
    12:30 beam back, left dipole set.
    12:35 current scan to study beam heating effect on He3 target.
    ased for 85 uA (90 on our BCM) then 10 uA step decrease to 15 uA.
    13:15 MCC inserted half-waveplate.
    13:26 run 1549 endrun failed. TS11 stuck. Rebooted and restarted CODA. All is fine. Need to repeat run with 75uA.
    BCM reading strange: when 65uA (MCC), it reads 118 uA (due to wrong calibration constants).
    14:22 changed BCM calibration constant from ~78 to ~45.
    14:52-14:55 beam off due to RF separator down.


    run 1544 7:49-9:48, 3.19M, 90 uA on He3, kin. #8.
    run 1545 9:49-10:01, 0.12M, 90 uA on He3. kin #8. Stopped for Mott measurement.

    Current scan:
    run 1546 12:32-12:42, 0.75M, 90 uA/(85uA MCC).
    run 1547, 12:46-12:59,0.95M, 84uA/78. Current varies since MCC doing adjustment.
    run 1548, 13:03-13:12,0.56M, 81uA/75. MCC still adjusting.
    run 1549, 13:16-13:26, 0.70M,80/75uA. ENDRUN failed, TS11 stuck.
    run 1550, 13:41-13:51, 0.56M, 118uA?/65uA. BCM reading problem.
    run 1551, 13:54-14:04,0.66M, 137uA?/75, BCM reading problem.
    run 1552, 14:11-14:21,0.55M, 101uA?/55, BCM reading problem.
    run 1553, 14:24-14:39,0.64M, 48uA/45. changed BCm calibration reading from ~78 to ~45 before run started.
    run 1554,14:42-14:54, 0.3M, 37uA/35. Beam not stable, RF separator trip.
    run 1555, 14:56-15:02,0.08M, 37uA/35, junk.
    run 1556, 15:03-15:18,0.46M, 37uA/35.
    run 1557, 15:20-15:32, 0.5M, 26uA/25.
    run 1558, 15:35-15:49, 0.39M, 16uA/15.
    run 1559, 15:51-?, 11uA/10.